write an equation in slope-intercept form of a line with the following characteristics: parallel to the graph of: x+y=6 and passes through (1,-4).
Answer by rfer(16322)

y=-x+6
-4=-m+6
m=10
----------------
y=mx+b
-4=10*1+b
b=-14
--------------
y=10x-14